A PbF PROBE FOR THE ELECTRON ELECTRIC DIPOLE MOMENT

\underline{JOHN MOORE-FURNEAUX}, N.E. SHAFER-RAY, Home L. Dodge Department of Physics and Astronomy, University of Oklahoma, Norman OK, 73019.

A molecule in a state of total angular momentum F has a familiar 2F + 1 degeneracy. In the presence of a pure magnetic field, this 2F + 1 degeneracy is completely lifted with each magnetic sub-level M_{F} acquiring its own energy. In the presence of an electric field, quantum states with non-zero magnetic quantum numbers |M_{F} | remain doubly degenerate. This fact is well established by Stark spectroscopy and is a consequence of time-reversal symmetry. In 1950, Purcell and Ramsey pointed out that time-reversal symmetry could be broken by the existence of an electric dipole moment of the electron. Over the last 60 years, the interest in such a symmetry breaking dipole moment has increased, in part because it may explain the matter-antimatter asymmetry of the Universe, and in part because it could help to differentiate between competing fundamental models of Physics. If large enough, such a dipole moment could lead to an observable lifting of the degeneracy between +M_{F} and −M_{F} states of a molecule in a pure electric field. We report on progress toward searching for an electric dipole moment by an optical quantum beat experiment utilizing the X_{1} ^{2}Π_{1/2} state of PbF molecule.